OpenStudy (anonymous):

the area of a rectangle is 108 sq feet . The length is 3ft more than the width. What are the length and width in feet ?

OpenStudy (anonymous):

(x+3)x=108 blah blah blah x = 9, x = -12

OpenStudy (anonymous):

Yeah, from the quadratic equation x^2 + 3x - 108 = 0 ... factorize to (x+9)(x-12) = 0, meaning that the two possible solutions are 9 and -12. Since you can't realistically have -12 feet, the answer must be 9ft for the original length. Therefor the width is 12.